### v2

# A linear Convergence result for the Jacobi-Proximal Alternating Direction Method of Multipliers

Abstract.
In this paper, we analyze the convergence rate of the Jacobi-Proximal Alternating Direction Method of Multipliers (ADMM), a method initially introduced by Deng et al.[6] for the block-structured optimization problem with linear constraint. We establish the linear convergence of the algorithm when the cost functions are strongly convex and smooth.
Key words and phrases:
Keywords: Alternating Direction Method of Multipliers, Jacobi-Proximal ADMM, Linear Convergence, Convex Optimization
1. Introduction
We consider the following convex optimization problem with $N$-block variables involving a linear constraint given as follows:
$\begin{split}\min_{x_{1},\cdots,x_{N}}&\sum_{i=1}^{N}\tilde{f}_{i}(x_{i})\\ \textrm{subject to}~{}&\sum_{i=1}^{N}A_{i}x_{i}=c,\\ &x_{i}\in\mathcal{X}_{i},\qquad i=1,\cdots,N.\end{split}$ (1.1)
where $N\geq 2$, $x_{i}\in\mathbb{R}^{n_{i}},~{}A_{i}\in\mathbb{R}^{m\times n_{i}},~{}c\in\mathbb{R}^{m}~{}\textrm{and~{}}\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ are closed convex sets and $\tilde{f}_{i}:\mathbb{R}^{n_{i}}\to\left(-\infty,+\infty\right]~{}(i=1,2,\cdots,N)$ are closed proper convex functions. The solution set of (1.1) is assumed to be nonempty. Note that the convex constraints $x_{i}\in\mathcal{X}_{i}$ can be incorporated into the objective $\tilde{f}_{i}$ using the indicator function:
$I_{\mathcal{X}_{i}}(x_{i})=\left\{\begin{array}[]{ll}0&\textrm{if }x_{i}\in\mathcal{X}_{i},\\ +\infty&\textrm{otherwise.}\end{array}\right.$ (1.2)
Then the problem (1.1) can be rewritten as
$\begin{split}\min_{x_{1},\cdots,x_{N}}&\sum_{i=1}^{N}f_{i}(x_{i})\\ \textrm{subject to}~{}&\sum_{i=1}^{N}A_{i}x_{i}=c,\end{split}$ (1.3)
where $f_{i}(x_{i})=\tilde{f}_{i}(x_{i})+I_{\mathcal{X}_{i}}(x_{i})$. Thus, we consider the equivalent reformulation (1.1) throughout this paper for ease of presentation.
The problem (1.1) appears in various applications that contain the image alignment problem in [33], the robust principal component analysis model with noisy and incomplete data in [36], the latent variable Gaussian graphical model selection in [3, 31] and the quadratic discriminant analysis model in [30].
To solve the problem (1.1), various algorithms have been proposed. A simple distributed algorithm for solving (1.1) is dual decomposition [11], which is essentially a dual ascent method or dual subgradient method [35] as follows. Consider the Lagrangian for problem (1.1):
$\mathcal{L}(x_{1},\cdots,x_{N},\lambda)=\sum_{i=1}^{N}f_{i}(x_{i})-\bigg{\langle}\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\bigg{\rangle}$ (1.4)
where $\lambda\in\mathbb{R}^{m}$ is the Lagrangian multiplier or the dual variable. The dual decomposition updates as follows: for $k\geq 1$,
$\left\{\begin{array}[]{ll}\left(x_{1}^{k+1},x_{2}^{k+1},\cdots,x_{N}^{k+1}\right)=\arg\min_{\{x_{i}\}}\mathcal{L}(x_{1},\cdots,x_{N},\lambda^{k}),\\ \lambda^{k+1}=\lambda^{k}-\alpha_{k}\left(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\right),\end{array}\right.$
where $\alpha_{k}>0$ is a step-size. Since all the $x_{i}$’s are separable in the Lagrangian function (1.4), the $x$-update step reduces to solving $N$ individual $x_{i}$- subproblems:
$x_{i}^{k+1}=\arg\min_{x_{i}}f_{i}(x_{i})-\big{\langle}\lambda^{k},A_{i}x_{i}\big{\rangle}\quad\textrm{for}~{}i=1,2,\cdots,N,$
and thus they can be carried out in parallel. With a suitable choice of $\alpha_{k}$ and certain assumptions, dual decomposition is guaranteed to converge to an optimal solution [35]. However, dual decomposition exhibits slow convergence in practice, with a convergence rate of $\mathcal{O}(1/\sqrt{k})$ for general convex problems.
An effective distributed algorithm for (1.1) can be designed based on the alternating direction method of multipliers (ADMM) [13, 15] using the following augmented Lagrangian:
$\mathcal{L}_{\rho}(x_{1},\cdots,x_{N};\lambda):=\sum_{i=1}^{N}f_{i}(x_{i})-\bigg{\langle}\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\bigg{\rangle}+\frac{\rho}{2}\bigg{\|}\sum_{i=1}^{N}A_{i}x_{i}-c\bigg{\|}^{2},$
where $\rho>0$ is a penalty parameter. In the context of ADMM using the augmented Lagrangian, the primal variable $x_{i}$ can be updated using different approaches. One common method is the Gauss-Seidel update, in which the primal variable $x_{i}$ is updated sequentially as follows:
$\begin{split}x_{i}^{k+1}&=\arg\min_{x_{i}}\mathcal{L}_{\rho}(x_{1}^{k+1},\cdots,x_{i-1}^{k+1},x_{i},x_{i+1}^{k},\cdots,x_{N}^{k},\lambda^{k})\\ &=\arg\min_{x_{i}}f_{i}(x_{i})+\frac{\rho}{2}\bigg{\|}\sum_{j<i}A_{j}x_{j}^{k+1}+A_{i}x_{i}+\sum_{j>i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\bigg{\|}_{2}^{2}.\end{split}$
Alternatively, the Jacobi-type update provides a parallel update scheme for the primal variable $x_{i}$ as follows:
$\begin{split}x_{i}^{k+1}&=\arg\min_{x_{i}}\mathcal{L}_{\rho}(x_{1}^{k},\cdots,x_{i-1}^{k},x_{i},x_{i+1}^{k},\cdots,x_{N}^{k},\lambda^{k})\\ &=\arg\min_{x_{i}}f_{i}(x_{i})+\frac{\rho}{2}\bigg{\|}A_{i}x_{i}+\sum_{j\neq i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\bigg{\|}_{2}^{2}.\end{split}$
Since Gauss-Seidel ADMM updates the primal variables $x_{i}$ using the most recent information, the performance of Jacobi ADMM is typically expected to be worse than that of Gauss-Seidel ADMM. For instance, the work [19] found an example for which Jacobi ADMM diverges even when $N=2$. This issue was addressed by Deng et al.[6] that introduced the Jacobi-Proximal ADMM involving a proximal term $\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}$ for each $x_{i}$-subproblem with positive semi-definite matrix $P_{i}$ and a damping parameter $\gamma>0$ for the update of $\lambda$ to guarantee its convergence. The Jacobi-Proximal ADMM is formulated as in Algorithm 1.
Algorithm 1 Jacobi-Proximal ADMM
  Initialize: $x_{i}^{0}~{}(i=1,2,\cdots,N)$ and $\lambda^{0}$;
  for $k=0,1,\cdots$ do
     Update $x_{i}$ for $i=1,\cdots,N$ in parallel by:
$x_{i}^{k+1}=\arg\min_{x_{i}}\left\{f_{i}(x_{i})+\frac{\rho}{2}\Bigg{\|}A_{i}x_{i}+\sum_{j\neq i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\Bigg{\|}_{2}^{2}+\frac{1}{2}\big{\|}x_{i}-x_{i}^{k}\big{\|}_{P_{i}}^{2}\right\};$ (1.5)
Update $\textrm{Update }\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg{(}\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg{)}.$ (1.6)
  end for
Here we denote $\|x_{i}\|^{2}_{P_{i}}:=x_{i}^{T}P_{i}x_{i}$. The Jacobi-Proximal ADMM has the advantage of being highly parallelizable, as the updates for $x_{i}$ for each $i=1,2,\ldots,N$ can be computed simultaneously for each $i$, offering potential computational efficiency gains. We note that the original $x_{i}$-subproblem (1.5) contains a quadratic term of the form $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$. If $A_{i}^{T}A_{i}$ is ill-conditioned or computationally expensive to invert, one approach is to modify $P_{i}$ as $P_{i}=D_{i}-\rho A_{i}^{T}A_{i}$, where $D_{i}$ is a well-conditioned and simple matrix(e.g., a diagonal matrix). This modification eliminates the quadratic term $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$, making the $x_{i}$-subproblem easier to solve. There are various choices for $P_{i}$ that can simplify the computation of the $x_{i}$-subproblem. Here are two widely used formulations for $P_{i}$ in [7], corresponding to the following methods:
  1. (1)
    The standard proximal method, where $P_{i}$ is taken as $P_{i}=\tau_{i}\mathbf{I}$, with $\tau_{i}>0$.
  2. (2)
    The prox-linear method, where $P_{i}$ is taken as $P_{i}=\tau_{i}\mathbf{I}-\rho A_{i}^{T}A_{i}$, with $\tau_{i}>0$.
The work of Deng et al.[6] established a sublinear convergence result for the convergence of Jacobi-Proximal ADMM when the proximal matrices $P_{i}$ satisfy $P_{i}\succ(\frac{N}{2-\gamma}-1)A_{i}^{T}A_{i}$ for $i=1,2,\ldots,N$ with $\gamma<2$. In particular, under specific choices of $P_{i}$, the convergence conditions reduce to $\tau_{i}>\rho(\frac{N}{2-\gamma}-1)\|A_{i}\|^{2}$ for the standard proximal method and $\tau_{i}>\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}$ for the prox-linear method.
We then extend the analysis to establish the linear convergence of Jacobi-Proximal ADMM. It is well-known that ADMM converges linearly when the cost functions exhibit additional structure, such as strong convexity. We prove that the Jacobi-Proximal ADMM converges linearly under the assumption that each cost function $f_{j}$ is strongly convex. We derive a condition on the parameters $\rho>0$ and $\gamma>0$ for linear convergence, and we establish a linear convergence rate dependent on these parameters. We refer to Section 2 for the details and we provide a comparison of our convergence result with previous work on the problem (1.1) in Table 1.
Ref. Convexity Regularity Rank Conditions Rate
Gauss-Seidel ADMM for $N=2$ [7] $f_{1}$: SC $f_{1}$: $L$-smooth $A_{1}$: full column rank $A_{2}$: full row rank linear
$f_{1},f_{2}$: SC $f_{1}$: $L$-smooth $A_{2}$: full row rank
$f_{1}$: SC $f_{1},f_{2}$: $L$-smooth $A_{1}$: full column rank
$f_{1},f_{2}$: SC $f_{1},f_{2}$: $L$-smooth
Gauss-Seidel ADMM for $N\geq 3$ [26] $f_{2},\ldots,f_{N}$: SC $f_{N}$: $L$-smooth $A_{N}$: full row rank linear
$f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth
$f_{2},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $A_{1}$: full column rank
Gauss-Sediel ADMM for $N\geq 3$ [22] $f_{i}(x_{i})=g_{i}(E_{i}x_{i})+h_{i}(x_{i})$ $g_{i}$: strictly C $h_{i}$: C $g_{i}$: $L$-smooth $E_{i}$: full column rank linear
Jacobi-Proximal ADMM for $N\geq 3$ [6] $f_{1},\ldots,f_{N}$: C $A_{i}$: full column rank sub-linear
Jacobi-Proximal ADMM for $N\geq 3$ This work $f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}$: full row rank linear
Table 1. This table summarizes the convergence results for ADMM. In the ’Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex). In the ’Rate’ column, ’linear’ refers to the cases with linear convergence, while ’sublinear’ indicates the case with sublinear convergence. $N$ denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
1.1. Related works
The Alternating Direction Method of Multipliers (ADMM) [13, 15] is an effective approach that simplifies the original problem by decomposing it into multiple easier-to-solve minimization subproblems. Building on the classical Douglas-Rachford operator splitting method [8], Glowinski and Marrocco [16], along with Gabay and Mercier [14], introduced the well-known alternating direction method of multipliers (ADMM) with Gauss-Seidel decomposition for solving (1.1) when $N=2$. It utilizes the augmented Lagrangian for (1.1):
$\mathcal{L}_{\rho}(x_{1},\cdots,x_{N};\lambda):=\sum_{i=1}^{N}f_{i}(x_{i})-\left<\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\right>+\frac{\rho}{2}\left\|\sum_{i=1}^{N}A_{i}x_{i}-c\right\|^{2}$
which incorporates a quadratic penalty of the constraints (with a parameter $\rho>0$) into the Lagrangian. At each iteration, the augmented Lagrangian is sequentially minimized with respect to $x_{1}$ and $x_{2}$, followed by an update of the dual variable $\lambda$. The ADMM has been extensively investigated for solving two-block convex minimization problems (i.e., $N=2$). Its global convergence in the case of $N=2$ has been established in [9, 12]. In a more recent work by Deng and Yin [7], a generalized version of ADMM was introduced, where proximal terms were added to the two subproblems in (2.2), demonstrating linear convergence under certain assumptions on the strong convexity of functions $f_{1}$ and $f_{2}$, as well as the rank conditions on $A_{1}$ and $A_{2}$. Additionally, various studies have explored the convergence rate properties of ADMM when $N=2$ (see, e.g., [28, 10, 18, 1, 29, 20, 32, 21]). Although the convergence properties of ADMM for $N=2$ are well understood, its behavior for $N\geq 3$ has remained uncertain for a considerable period. In the work by Chen et al. [4], a counterexample was provided, demonstrating that without additional conditions, the ADMM for $N\geq 3$ may fail to converge. Building on this, we first discuss studies that establishes the convergence conditions for ADMM using the Gauss-Seidel scheme when $N\geq 3$, where the Gauss-Seidel method updates variables sequentially. We refer to the works [2], [23], [5], and [27] that achieved global and sublinear convergence results of the 3-block ADMM. Han and Yuan [17] established the global convergence of ADMM (2.2) under the assumption that $f_{1},\cdots,f_{N}$ are all strongly convex and $\gamma$ is restricted to a specific range. Both Chen, Shen, and You [5] and Lin, Ma, and Zhang [25] relax the conditions compared to [17] by allowing only $N-1$ functions to be strongly convex, with with Chen et al. [5] considering the case $N=3$ and Lin et al. [25] considering $N\geq 3$. Furthermore, Lin, Ma, and Zhang [25] demonstrated that under these conditions, ADMM (2.2) exhibits a sublinear convergence rate in both the ergodic and non-ergodic senses. Lin, Ma, and Zhang [24] further presented sufficient conditions to guarantee a sublinear convergence rate for ADMM when $N\geq 3$. Lin, Ma, and Zhang [26] proved the global linear convergence of ADMM under the three scenarios. The global R-linear convergence of ADMM has been established by Hong and Luo [22]. Alternatively, Jacobi updates in ADMM have been extensively studied, offering parallelization advantages for multi-block optimization problems. The linear convergence of Gauss-Seidel ADMM, Proximal ADMM, and modified Jacobi ADMM, as shown in [22], is established under several assumptions. In [6], it was established that the multi-block Jacobi ADMM converges globally at a rate of $o(1/k)$ in two cases: (i) when all coefficient matrices are nearly mutually orthogonal and possess full column rank, or (ii) when proximal terms are introduced into the subproblems. In [19], the multi-block Jacobi ADMM was combined with a relaxation step, achieving a worst-case convergence rate of $O(1/k)$ in both ergodic and non-ergodic senses. In a recent work by Xiao, Deng, and Xu [38], the authors introduce the Distributed Jacobi-Proximal ADMM algorithm for solving consensus optimization problems.
1.2. Contribution
The main contribution in this paper are as follows. We establish the global linear convergence of Jacobi-Proximal ADMM(Algorithm 1) under the assumption that the objective function $f_{i}$ are strictly convex and $L_{i}$-smooth. We establish its convergence at a rate of $O(\sigma^{k})$ for the optimality measure defined in (2.4). Our numerical results on the LCQP model and the optimal resource allocation problem support the linear convergence result of this work.
1.3. Organization
We organize the rest of this paper as follows. In Section 2, we present a sufficient condition to ensure the convergence of (2.2). Then based on the analysis in Section 2, we prove the global convergence of the Algorithm 1 in Section 3. In Section 4, we present convergence results by conducting numerical experiments on two examples. Finally, we conclude the paper in Section 5.
2. Statement of the main result
In this section, we state the main result of this paper concerning the linear convergence of the Jacobi-Proximal ADMM in Algorithm 1. Since both of the objective function and constraints of (1.3) are sums of terms on individual $x_{i}$’s, the problem can be decomposed into $N$ smaller subproblems, which is solved in a parallel and distributed manner.
Related to the problem (1.3), we first define the following augmented Lagrangian
$\mathcal{L}_{\rho}(x_{1},\cdots,x_{N};\lambda):=\sum_{i=1}^{N}f_{i}(x_{i})-\left<\lambda,\sum_{i=1}^{N}A_{i}x_{i}-b\right>+\frac{\rho}{2}\left\|\sum_{i=1}^{N}A_{i}x_{i}-b\right\|^{2}$
where $\lambda\in\mathbb{R}^{m}$ is the dual variable and parameter $\rho>0$.
We state assumptions on the cost functions which will be used throughout the paper. Here we let $n=n_{1}+\cdots+n_{N}$ and set $A\in\mathbb{R}^{m\times n}$ by $A=\big{[}A_{1}\cdots A_{N}\big{]}$. Also, we set $x^{k}\in\mathbb{R}^{n}$ for $k\geq 0$ and $x^{*}\in\mathbb{R}^{n}$ as $x^{k}=\big{[}x_{1}^{k};~{}\cdots;~{}x_{N}^{k}\big{]}$ and $x^{*}=\big{[}x_{1}^{*};~{}\cdots;~{}x_{N}^{*}\big{]}$. We then have
$Ax^{k}=\sum_{i=1}^{N}A_{i}x_{i}^{k}.$
Assumption 2.1.
The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\left(-\infty,+\infty\right]$ are closed proper convex for $1\leq i\leq N$.
Assumption 2.2.
We assume that $x^{*}=(x_{1}^{*},\cdots,x_{N}^{*})\in\mathbb{R}^{n}$ is a minimizer of problem (1.3) companied with a multiplier $\lambda^{*}\in\mathbb{R}^{m}$ satisfying the following KKT condition
$\nabla f_{i}\left(x_{i}^{*}\right)=A_{i}^{T}\lambda^{*}\quad\textrm{for}~{}i=1,\cdots,N,$ (2.1)
$Ax^{*}=\sum_{i=1}^{N}A_{i}x_{i}^{*}=c.$ (2.2)
Assumption 2.3.
There are positive constants $L_{i}$ such that for all $x,y\in\mathbf{R}^{n_{i}}$,
$\left\|\nabla f_{i}(x)-\nabla f_{i}(y)\right\|\leq L_{i}\left\|x-y\right\|\textrm{ for }1\leq i\leq N.$
Assumption 2.4.
The function $f_{i}:\mathbb{R}^{n_{i}}\to(-\infty,+\infty]$ is $\alpha$-strongly convex for some $\alpha>0$, $i.e.$,
$f_{i}(y)\geq f_{i}(x)+\left\langle y-x,\nabla f_{i}(x)\right\rangle+\alpha\left\|x-y\right\|^{2}$
for all $x,y\in\mathbb{R}^{n_{i}}$.
Assumption 2.5.
We assume that the matrix $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}\in\mathbb{R}^{n\times m}$ with $n=n_{1}+\cdots+n_{N}$ has full column rank.
Here $P_{i}$ is positive semi-definite for each $1\leq i\leq N$. For $k\geq 0$, we denote by $\left(x_{1}^{k},\cdots,x_{N}^{k};\lambda^{k}\right)$ the variable generated by the Jacobi-Proximal ADMM. We state the main result of this paper.
Theorem 2.6.
Suppose Assumptions 2.1-2.5 hold. Choose any $s>0$ such that
$s\Big{[}\rho^{2}D\|A_{i}\|+\frac{L}{N}\Big{]}<\frac{\alpha}{2N},\qquad\forall 1\leq i\leq N,$
and suppose that $20<\gamma<$ 2, $\rho>0$ are chosen so that there exist values $\xi_{i}>0$ such that
$\left\{\begin{array}[]{ll}\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}\succ 0,\\ \sum_{i=1}^{N}\xi_{i}<2-\gamma.\end{array}\right.$ (2.3)
Let $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}$. Here $c_{A}>0$ and $\mu_{s}\in(0,1)$ are constants satisfying
[…]
where the first inequality ensures that $c_{A}=\sqrt{\min\textrm{eig}\Big{(}\sum_{i=1}^{N}A_{i}A_{i}^{T}\Big{)}}>0$. Then we have
$\phi\big{(}x^{k+1},\lambda^{k+1}\big{)}\leq\sigma\phi\big{(}x^{k},\lambda^{k}\big{)},$ (2.4)
where $\phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is defined as $\phi\big{(}x^{k},\lambda^{k}\big{)}=\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}$. As a result of (2.4), we have the following result
$\phi\big{(}x^{k},\lambda^{k}\big{)}=O(\sigma^{k}).$
Remark 2.7.
Choosing $\xi_{i}<\frac{2-\gamma}{N}$ leads to a simplification of the convergence condition (2.3), resulting in the following form:
$\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho N}{2-\gamma}A_{i}^{T}A_{i}\succ 0.$ (2.5)
In the case of $P_{i}=\tau_{i}\mathbf{I}$ (standard proximal method), the condition (2.5) further reduces to
$\rho\|A_{i}\|^{2}+\tau_{i}-8s(\rho\|A_{i}\|^{2}+\tau_{i})^{2}-\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}>0.$
Similarly, for $P_{i}=\tau_{i}\mathbf{I}-\rho A_{i}^{T}A_{i}$ (prox-linear method), the condition (2.5) simplifies to
$\tau_{i}-8s\tau_{i}^{2}-\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}>0.$
3. Convergence
In this section, we prove Theorem 2.6. For this, we prepare the several lemmas.
Lemma 3.1.
Suppose Assumption 2.5 holds. Consider matrices $A_{i}\in\mathbb{R}^{m\times n_{i}}$ for $1\leq i\leq N$. Then there exists $c_{A}>0$ such that
$\bigg{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\bigg{)}^{\frac{1}{2}}\geq c_{A}\|\lambda\|$ (3.1)
for all $\lambda\in\mathbb{R}^{m}$.
Proof.
We define $c_{A}\geq 0$ by
$c_{A}=\min_{\lambda\in\mathbb{R}^{m}\backslash\{0\}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}=\min_{\lambda\in S^{m-1}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}.$
In order to prove that $c_{A}>0$, we assume $c_{A}=0$ and proceed to derive a contradiction. Since $S^{m-1}$ is compact, there exists a minimizer $\lambda_{0}\in S^{m-1}$ of (3.1) such that
$\frac{\Big{(}\sum_{i=1}^{N}\big{\|}A_{i}^{T}\lambda_{0}\big{\|}^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda_{0}\|}=0.$
Thus we have $A_{i}^{T}\lambda_{0}=0$ for all $1\leq i\leq N$, and this leads to $\lambda_{0}=0$ since $\big{[}A_{1}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}$ has full column rank. It contradicts to the fact that $\lambda_{0}\in S^{m-1}$. Therefore, we have $c_{A}>0$, and the lemma is proved. ∎
Remark 3.2.
The assumption of Lemma 3.1 means that the matrix $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}\in\mathbb{R}^{n\times m}$ with $n=n_{1}+\cdots+n_{N}$ has only a trivial kernel, i.e., $\ker\left(\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}\right)=\{\mathbf{0}\}$.
By the strongly convexity of the objective function, we have the following lemma.
Lemma 3.3.
Suppose Assumptions 2.1, 2.2, and 2.4 hold. Then it holds that
$\begin{split}&\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\geq\alpha\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+H_{k}\end{split}$ (3.2)
where $H_{k}:=\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}$.
Proof.
By the optimality of (1.5), we have
$\nabla f_{i}(x_{i}^{k+1})=A_{i}^{T}\lambda^{k}-\rho A_{i}^{T}\bigg{(}A_{i}x_{i}^{k+1}+\sum_{j\neq i}A_{j}x_{j}^{k}-c\bigg{)}+P_{i}\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}.$ (3.3)
Combining (3.3) and (2.1), we get
$\begin{split}&A_{i}^{T}\big{(}\lambda^{k}-\lambda^{*}\big{)}\\ &=\nabla f_{i}\big{(}x_{i}^{k+1}\big{)}-\nabla f_{i}\left(x_{i}^{*}\right)+\rho A_{i}^{T}\bigg{(}A_{i}x_{i}^{k+1}+\sum_{j\neq i}A_{j}x_{j}^{k}-c\bigg{)}+P_{i}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}.\end{split}$ (3.4)
Since each $f_{i}$ is $\alpha$-strongly convex, we have
$\left\langle x_{i}^{k+1}-x_{i}^{*},\nabla f_{i}\big{(}x_{i}^{k+1}\big{)}-\nabla f_{i}\big{(}x_{i}^{*}\big{)}\right\rangle\geq\alpha\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}.$ (3.5)
Inserting (3.4) into (3.5), we have
$\begin{split}\bigg{\langle}A_{i}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)},\lambda^{k}-\lambda^{*}-\rho\bigg{(}A_{i}x_{i}^{k+1}+\sum_{j\neq i}A_{j}x_{j}^{k}-c\bigg{)}\bigg{\rangle}\\ +\left\langle x_{i}^{k+1}-x_{i}^{*},P_{i}\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\right\rangle\geq\alpha\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}.\end{split}$
We rewrite above inequality as follows
$\begin{split}\bigg{\langle}A_{i}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)},\lambda^{k}-\lambda^{*}-\rho\bigg{(}\sum_{i=1}^{N}A_{i}x_{i}^{k+1}+\sum_{j\neq i}A_{j}\big{(}x_{j}^{k}-x_{j}^{k+1}\big{)}-c\bigg{)}\bigg{\rangle}\\ +\left\langle x_{i}^{k+1}-x_{i}^{*},P_{i}\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\right\rangle\geq\alpha\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}.\end{split}$
For simplicity, we let $\hat{\lambda}=\lambda^{k}-\rho\left(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\right)=\lambda^{k}-\frac{1}{\gamma}\left(\lambda^{k}-\lambda^{k+1}\right)$. Then the above inequality reads as follows
$\begin{split}\bigg{\langle}A_{i}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)},\hat{\lambda}-\lambda^{*}-\rho\Big{(}\sum_{j\neq i}A_{j}(x_{j}^{k}-x_{j}^{k+1})\Big{)}\bigg{\rangle}\\ +\left\langle x_{i}^{k+1}-x_{i}^{*},P_{i}\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\right\rangle\geq\alpha\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}.\end{split}$
Using that $\sum_{j\neq i}A_{j}\big{(}x_{j}^{k}-x_{j}^{k+1}\big{)}=A\big{(}x^{k}-x^{k+1}\big{)}-A_{i}\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}$ and summing the above estimate for $1\leq i\leq N$ yields
$\begin{split}\left\langle A\big{(}x^{k+1}-x^{*}\big{)},\hat{\lambda}-\lambda^{*}\right\rangle+\sum_{i=1}^{N}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)}^{T}\left(\rho A_{i}^{T}A_{i}+P_{i}\right)\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\\ \geq\alpha\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\rho\Big{\langle}A\big{(}x^{k+1}-x^{*}\big{)},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}.\end{split}$ (3.6)
Combining the update rule (1.6) and (2.2) and using that $\hat{\lambda}-\lambda^{*}=\left(\lambda^{k+1}-\lambda^{*}\right)+\frac{\gamma-1}{\gamma}\left(\lambda^{k}-\lambda^{k+1}\right)$ in the above inequality, we get
$\begin{split}&\left\langle\frac{1}{\gamma\rho}\big{(}\lambda^{k}-\lambda^{k+1}\big{)},\lambda^{k+1}-\lambda^{*}+\frac{\gamma-1}{\gamma}\big{(}\lambda^{k}-\lambda^{k+1}\big{)}\right\rangle\\ &+\sum_{i=1}^{N}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)}^{T}\left(\rho A_{i}^{T}A_{i}+P_{i}\right)\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\\ &\geq\alpha\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\frac{1}{\gamma}\left\langle\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\right\rangle.\end{split}$ (3.7)
For the two terms in the left hand side, we have the following identities
$\begin{split}&\left\langle\frac{1}{\gamma\rho}\big{(}\lambda^{k}-\lambda^{k+1}\big{)},\lambda^{k+1}-\lambda^{*}+\frac{\gamma-1}{\gamma}\big{(}\lambda^{k}-\lambda^{k+1}\big{)}\right\rangle\\ &=\frac{\gamma-2}{2\gamma^{2}\rho}\left\|\lambda^{k+1}-\lambda^{k}\right\|^{2}-\frac{1}{2\gamma\rho}\left\|\lambda^{k+1}-\lambda^{*}\right\|^{2}+\frac{1}{2\gamma\rho}\left\|\lambda^{k}-\lambda^{*}\right\|^{2},\end{split}$
and
$\begin{split}&\sum_{i=1}^{N}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)}^{T}\left(\rho A_{i}^{T}A_{i}+P_{i}\right)\big{(}x_{i}^{k}-x_{i}^{k+1}\big{)}\\ &=-\frac{1}{2}\sum_{i=1}^{N}\Big{(}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\Big{)}.\end{split}$
Inserting the above two equalities in (3.7), we obtain
$\begin{split}&\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\geq\alpha\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+H_{k}\end{split}$
where $H_{k}=\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}$. This is the desired inequality. ∎
We notice that the term $\alpha\sum_{i=1}^{N}\left\|x_{i}^{k+1}-x_{i}^{*}\right\|^{2}$ in the inequality (3.2) is a gain for the strongly convexity of the costs. For the linear convergence, we make it to bound the distance $\left\|\lambda^{k}-\lambda^{*}\right\|^{2}$ between the multipliers $\lambda^{k}$ and $\lambda^{*}$ in the following lemma.
Lemma 3.4.
Suppose Assumption 2.1-2.5 hold. Let $L=\max_{1\leq i\leq N}L_{i}^{2}$ and $D=\sum_{i=1}^{N}\left\|A_{i}^{T}\right\|^{2}$. Then it holds that
$\begin{split}&c_{A}^{2}\left\|\lambda^{k}-\lambda^{*}\right\|^{2}\\ &\leq 2L\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+4\rho^{2}D\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}\\ &\quad+4\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}.\end{split}$ (3.8)
Proof.
Using Lemma 3.1 and (3.4), we get
$\begin{split}c_{A}^{2}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}&\leq\sum_{i=1}^{N}\big{\|}A_{i}^{T}(\lambda^{k}-\lambda^{*})\big{\|}^{2}\\ &\leq\sum_{i=1}^{N}\bigg{\|}\nabla f_{i}\big{(}x_{i}^{k+1}\big{)}-\nabla f_{i}\left(x_{i}^{*}\right)+\rho A_{i}^{T}\Big{(}A_{i}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)}+\sum_{j\neq i}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\Big{)}\\ &\qquad+P_{i}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}.\end{split}$
We apply the Cauchy-Schwartz inequality to deduce
$\begin{split}&c_{A}^{2}\left\|\lambda^{k}-\lambda^{*}\right\|^{2}\\ &\leq 2\sum_{i=1}^{N}\Big{\|}\nabla f_{i}\big{(}x_{i}^{k+1}\big{)}-\nabla f_{i}\left(x_{i}^{*}\right)\Big{\|}^{2}\\ &\qquad+2\sum_{i=1}^{N}\bigg{\|}\rho A_{i}^{T}\Big{(}A_{i}\big{(}x_{i}^{k+1}-x_{i}^{*}\big{)}+\sum_{j\neq i}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\Big{)}+P_{i}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &=2\sum_{i=1}^{N}\Big{\|}\nabla f_{i}\big{(}x_{i}^{k+1}\big{)}-\nabla f_{i}\left(x_{i}^{*}\right)\Big{\|}^{2}\\ &\qquad+2\sum_{i=1}^{N}\bigg{\|}\rho A_{i}^{T}\Big{(}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\Big{)}+\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2},\end{split}$
where the latter equality holds by a simple rearrangement. Now we apply Assumption 2.3 and the Cauch-Schwartz inequality to obtain
$\begin{split}&c_{A}^{2}\left\|\lambda^{k}-\lambda^{*}\right\|^{2}\\ &\leq 2\sum_{i=1}^{N}L_{i}^{2}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}\\ &\qquad+4\rho^{2}\sum_{i=1}^{N}\bigg{\|}A_{i}^{T}\Big{(}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\Big{)}\bigg{\|}^{2}+4\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}.\end{split}$
Applying the Cauchy-Schwartz inequality once more, we obtain
$\begin{split}&c_{A}^{2}\left\|\lambda^{k}-\lambda^{*}\right\|^{2}\\ &\leq 2\left(\max_{1\leq i\leq N}L_{i}^{2}\right)\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}\\ &\qquad+4\rho^{2}\sum_{i=1}^{N}\big{\|}A_{i}^{T}\big{\|}^{2}\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}+4\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &=2L\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+4\rho^{2}D\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}\\ &\qquad+4\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}.\end{split}$
This proves the lemma. ∎
We derive the following result using Lemma 3.3 and Lemma 3.4.
Proposition 3.5.
Suppose Assumptions 2.1-2.5 hold. For $s>0$, we have the following inequality
$\begin{split}&\left(\frac{1}{2\gamma\rho}-sc_{A}^{2}\right)\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\geq\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}+G_{k},\end{split}$ (3.9)
where $H_{k}:=\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}$ and $G_{k}:=H_{k}-4s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}$.
Proof.
We recall from Lemma 3.3 the following inequality
$\begin{split}&\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\geq\alpha\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\\ &+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}.\end{split}$ (3.10)
Next we recall the inequality of Lemma 3.4 given as
$\begin{split}&c_{A}^{2}\left\|\lambda^{k}-\lambda^{*}\right\|^{2}\\ &\leq 2L\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}\\ &\quad+4\rho^{2}D\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}+4\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}.\end{split}$ (3.11)
Summing (3.11) multiplied by $s>0$ and (3.10), we get
$\begin{split}&\left(\frac{1}{2\gamma\rho}-sc_{A}^{2}\right)\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}+4s\rho^{2}D\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}\\ &\geq(\alpha-2Ls)\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}^{2}+\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\quad+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-4s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &\quad+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}.\end{split}$ (3.12)
We use the Cauchy-Schwartz inequality to find
$4s\rho^{2}D\bigg{\|}\sum_{j=1}^{N}A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\bigg{\|}^{2}\leq 4Ns\rho^{2}D\sum_{j=1}^{N}\|A_{j}\big{(}x_{j}^{k}-x_{j}^{*}\big{)}\|^{2}.$
Using this inequality in (3.12) and rearranging, we get
$\begin{split}&\left(\frac{1}{2\gamma\rho}-sc_{A}^{2}\right)\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}}^{2}\\ &\geq\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}\\ &\quad+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-4s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &\quad+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle},\end{split}$
which is the desired inequality. ∎
To show the main convergence result Proposition 2.6, we need the following two lemmas.
Lemma 3.6.
If $s>0$ satisfies
$s\Big{[}\rho^{2}D\|A_{i}\|^{2}+\frac{L}{N}\Big{]}<\frac{\alpha}{2N},$ (3.13)
there exists $\mu_{s}\in(0,1)$ such that
$\|x\|^{2}_{(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}}\leq\mu_{s}\|x\|^{2}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}\quad\forall~{}x\in\mathbb{R}^{n_{i}}.$ (3.14)
Proof.
If we choose $s>0$ that satisfies (3.13), we have
$(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}\prec\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}.$
Thus, we obtain the final result for $\mu_{s}\in(0,1)$, which is sufficiently close to $1$. ∎
Lemma 3.7.
If one chooses some $\xi_{i}>0$ such that $20<\gamma<$ 2, $\rho>0$, and $s>0$ satisfies the following conditions:
$\left\{\begin{array}[]{ll}\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}\succ 0,\\ \sum_{i=1}^{N}\xi_{i}<2-\gamma,\end{array}\right.$
then we have
$\begin{split}&\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-4s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &\quad+\frac{2-\gamma}{2\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{1}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}\geq 0.\end{split}$
Proof.
Using Young’s inequality, we have
$\begin{split}\frac{2}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}&=\sum_{i=1}^{N}\frac{2}{\gamma}(\lambda^{k}-\lambda^{k+1})^{T}A_{i}(x_{i}^{k}-x_{i}^{k+1})\\ &\geq-\sum_{i=1}^{N}\bigg{(}\frac{\xi_{i}}{\rho\gamma^{2}}\|\lambda^{k+1}-\lambda^{k}\|^{2}+\frac{\rho}{\xi_{i}}\|A_{i}(x_{i}^{k}-x_{i}^{k+1})\|^{2}\bigg{)}\\ &=-\frac{\sum_{i=1}^{N}\xi_{i}}{\rho\gamma^{2}}\|\lambda^{k+1}-\lambda^{k}\|^{2}-\sum_{i=1}^{N}\frac{\rho}{\xi_{i}}\|A_{i}(x_{i}^{k}-x_{i}^{k+1})\|^{2},\end{split}$ (3.15)
for $\xi_{j}>0~{}(1\leq j\leq N)$. Adding $\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}$, $-8s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}$, and $\frac{2-\gamma}{\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}$ to both sides of (3.15), we obtain
$\begin{split}&\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-8s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}\\ &\quad+\frac{2-\gamma}{\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}+\frac{2}{\gamma}\Big{\langle}\lambda^{k}-\lambda^{k+1},A\big{(}x^{k}-x^{k+1}\big{)}\Big{\rangle}\\ &\geq\sum_{i=1}^{N}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}_{\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}}\\ &+\bigg{(}\frac{2-\gamma-\sum_{i=1}^{N}\xi_{i}}{\gamma^{2}\rho}\bigg{)}\|\lambda^{k+1}-\lambda^{k}\|^{2}\end{split}$
for $\xi_{j}>0~{}(j=1,\cdots,N)$. The proof is done. ∎
We are now ready to show our main convergence and rate results.
Proof of Theorem 2.6.
By applying Lemma 3.6 and Lemma 3.7 to the inequality of Proposition 3.5, we get
$\begin{split}&\left(\frac{1}{2\gamma\rho}-sc_{A}^{2}\right)\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{\mu_{s}}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}\\ &\geq\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}.\end{split}$
This gives the following inequality
$\begin{split}&\sigma\bigg{(}\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}\bigg{)}\\ &\geq\frac{1}{2\gamma\rho}\big{\|}\lambda^{k+1}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2},\end{split}$
where $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}$. The proof is done. ∎
4. Numerical experiments
In this section, we present numerical tests of Algorithm 1 supporting the linear convergence property proved in Theorem 2.6. We test the Algorithm 1 for the linearly const rained quadratic programming (LCQP) model [34] and the optimal source allocation problem [37].
4.1. The LCQP Model
Here we consider the following LCQP model:
$\begin{split}&\min\sum_{i=1}^{N}f_{i}(x_{i})\\ &\textrm{s.t.}~{}\sum_{i=1}^{N}A_{i}x_{i}=c\end{split}$ (4.1)
where $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}~{}\textrm{is given by}~{}f_{i}(x_{i})=\frac{1}{2}x_{i}^{T}H_{i}x_{i}+x_{i}^{T}q_{i}\textrm{ with}~{}H_{i}\in\mathbb{S}_{+}^{n_{i}},~{}q_{i}\in\mathbb{R}^{n_{i}},~{}A_{i}\in\mathbb{R}^{m\times n_{i}}~{}\textrm{and}~{}c\in\mathbb{R}^{m}$. Since $H_{i}\in\mathbb{S}_{+}^{n_{i}}$, the quadratic function $f_{i}$ is strongly convex for each $i=1,\ldots,N$. We take $n_{1}=\cdots=n_{N}:=n$ in our experiment. We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1. The matrices involved in this problem are randomly generated through the following procedure:
  1. (1)
    Each $A_{i}\in\mathbb{R}^{m\times n_{i}}(i=1,\ldots,N)$ is randomly generated with i.i.d. standard normal entries with satisfying the condition that $\big{[}A_{1}^{T};\ldots;A_{N}^{T}\big{]}$ has full column rank.
  2. (2)
    The positive semi-definite matrices $P_{i}\in\mathbb{R}^{n_{i}\times n_{i}}(i=1,\ldots,N)$ are generated by sampling $B\in\mathbb{R}^{n_{i}\times n_{i}}$ with i.i.d. standard normal entries, and then symmetrizing as $P_{i}=\frac{1}{2}(B+B^{T})$ with checking that the eigenvalues of $P_{i}$ are non-negative.
  3. (3)
    The positive definite matrices $H_{i}\in\mathbb{S}_{+}^{n_{i}}(i=1,\ldots,N)$ are generated by first sampling $B\in\mathbb{R}^{n_{i}\times n_{i}}$ with i.i.d. standard normal entries, and then symmetrizing as $H_{i}=\frac{1}{2}(B+B^{T})$ with checking that the eigenvalues of $H_{i}$ are positive.
  4. (4)
    We first generate the optimal solution $x_{i}^{*}\in\mathbb{R}^{n_{i}}$ and $\lambda^{*}\in\mathbb{R}^{m}$ as follows: $x_{i}^{*}=randn(n_{i},1)$ for $i=1,\ldots,N$, and $\lambda^{*}=randn(m,1)$;
  5. (5)
    Next we define $q_{i}\in\mathbb{R}^{n_{i}}(i=1,\cdots,N)$ and $c\in\mathbb{R}^{m}$ as follows: $q_{i}=-H_{i}x_{i}^{*}+A_{i}^{T}\lambda^{*}$ and $c=\sum_{i=1}^{N}A_{i}x_{i}^{*}$ so that $\left({x_{1}^{*}}^{T},\cdots,{x_{N}^{*}}^{T},{\lambda^{*}}^{T}\right)^{T}$ satisfies the KKT conditions.
We update $x_{i}^{k+1}$ in (1.5) using the equation from the optimality given as follows:
$H_{i}x_{i}^{k+1}+q_{i}+\rho A_{i}^{T}\Bigg{(}A_{i}x_{i}^{k+1}+\sum_{j\neq i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\Bigg{)}+P_{i}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}=0.$
We define the error metric at each $k$-th iteration for $u^{k}=\big{(}x_{1}^{k},\ldots,x_{N}^{k};\lambda^{k}\big{)}$ as follows:
$\text{dis}(u^{k}):=\max\Big{\{}\big{\|}x_{1}^{k}-x_{1}^{*}\big{\|},\ldots,\big{\|}x_{N}^{k}-x_{N}^{*}\big{\|},\big{\|}\lambda^{k}-\lambda^{*}\big{\|}\Big{\}}.$ (4.2)
The initial values $u^{0}=(x_{1}^{0},\ldots,x_{N}^{0};\lambda^{0})$ are set to zero. The experiments were conducted for $N=3$ with $m=100$ and $n=40$, and for $N=10$ with $m=100$ and $n=60$. The experimental results are plotted in Figures 1 and 2 for the $3$-block case, and in Figures 3 and 4 for the $10$-block case. We find that the value $\text{dis}(u^{k})$ decreases to zero linearly as expected in Theorem 2.6.
Figure 1. Experimental results on 3-block LCQP model for fixed $\gamma$ under varying $\rho$.
Figure 2. Experimental results on 3-block LCQP model for fixed $\rho$ under varying $\gamma$.
Figure 3. Experimental results on 10-block LCQP model for fixed $\gamma$ under varying $\rho$.
Figure 4. Experimental results on 10-block LCQP model for fixed $\rho$ under varying $\gamma$.
4.2. The Optimal Resource Allocation Problem
In this example, we consider the following case:
$\begin{split}&\min\sum_{i=1}^{N}f_{i}(x_{i})\\ &\textrm{s.t. }\sum_{i=1}^{N}x_{i}=0\end{split}$
where $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is given as $f_{i}(x_{i})=(1/2)a_{i}(x_{i}-c_{i})^{2}+\log[1+exp(b_{i}(x_{i}-d_{i}))]$ with the coefficients $a_{i},b_{i},c_{i},d_{i}$ generated randomly with uniform distributions on $[0,2]$, $[-2,2]$, $[-10,10]$, $[-10,10]$ respectively. We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1. We update $x_{i}^{k+1}$ in (1.5) using the equation derived from the optimality given as follows:
$a_{i}\big{(}x_{i}^{k+1}-c_{i}\big{)}+\frac{b_{i}exp\Big{[}b_{i}\big{(}x_{i}^{k+1}-d_{i}\big{)}\Big{]}}{1+exp\left[b_{i}\big{(}x_{i}^{k+1}-d_{i}\big{)}\right]}+\rho\bigg{(}x_{i}^{k+1}+\sum_{j\neq i}x_{j}^{k}-\frac{\lambda^{k}}{\rho}\bigg{)}+P_{i}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}=0.$
In order to study the convergence behavior of the test algorithms, we test two cases: $6$-block and $20$-block. The value $u^{*}=\big{(}x_{1}^{*},\ldots,x_{N}^{*};\lambda^{*}\big{)}$ is obtained numerically by running Algorithm 1 $5,000$ iterations. We employ the same error metric $\text{dis}(u^{k})$ as defined in (4.2) for this experiment. The initial values $u^{0}=(x_{1}^{0},\ldots,x_{N}^{0};\lambda^{0})$ are set to zero. The experimental results are plotted in Figures 5 and 6 for the $6$-block case, and in Figures 7 and 8 for the $20$-block case. We observe that the value $\text{dis}(u^{k})$ converges linearly to zero, as anticipated in Theorem 2.6.
Figure 5. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
Figure 6. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
Figure 7. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
Figure 8. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
5. Conclusion
In this paper, we consider a multi-block Jacobi-Proximal ADMM algorithm for solving linearly coupled optimization problems in a parallel way. We established the linear convergence of the algorithm and derived its convergence rate involving more than two variable blocks when the cost functions are strongly convex and Lipschitz continuous. This is the first linear convergence result for the Jacobi-Proximal ADMM.

### v4

@@
 A linear convergence result for the Jacobi-Proximal Alternating Direction Method of Multipliers
-Abstract.
-In this paper, we analyze the convergence rate of the Jacobi-Proximal Alternating Direction Method of Multipliers (ADMM), a method initially introduced by Deng et al.[6] for the block-structured optimization problem with linear constraint.
+Abstract
+This paper is concerned with the Jacobi-Proximal Alternating Direction Method of Multipliers (Jacobi-Proximal ADMM) proposed in the works He, Xu, and Yuan (J.
+Sci.
+Comput., 2015) [On_the_proximal] and Deng et al.
+(J.
+Sci.
+Comput., 2017) [Parallel_multi-block], a parallel implementation of the ADMM for large-scale multi-block optimization problems with a linear constraint.
+While the $\mathit{o}(1/k)$ convergence of the Jacobi-Proximal ADMM for the case $N\geq 3$ has been well established in previous work, to the best of our knowledge, its linear convergence for $N\geq 3$ remains unproven.
 We establish the linear convergence of the algorithm when the cost functions are strongly convex and smooth.
-Key words and phrases:
-Keywords: Alternating Direction Method of Multipliers, Jacobi-Proximal ADMM, Linear Convergence, Convex Optimization
-1. Introduction
+Numerical experiments are presented supporting the convergence result.
+keywords:
+Alternating Direction Method of Multipliers, Jacobi-Proximal ADMM, Linear Convergence, Convex Optimization
+pacs:
+MSC Classification]90C25
+1 Introduction
 We consider the following convex optimization problem with $N$-block variables involving a linear constraint given as follows:
-$\begin{split}\min_{x_{1},\cdots,x_{N}}&\sum_{i=1}^{N}\tilde{f}_{i}(x_{i})\\ \textrm{subject to}~{}&\sum_{i=1}^{N}A_{i}x_{i}=c,\\ &x_{i}\in\mathcal{X}_{i},\qquad i=1,\cdots,N.\end{split}$ (1.1)
-where $N\geq 2$, $x_{i}\in\mathbb{R}^{n_{i}},~{}A_{i}\in\mathbb{R}^{m\times n_{i}},~{}c\in\mathbb{R}^{m}~{}\textrm{and~{}}\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ are closed convex sets and $\tilde{f}_{i}:\mathbb{R}^{n_{i}}\to\left(-\infty,+\infty\right]~{}(i=1,2,\cdots,N)$ are closed proper convex functions.
-The solution set of (1.1) is assumed to be nonempty.
-Note that the convex constraints $x_{i}\in\mathcal{X}_{i}$ can be incorporated into the objective $\tilde{f}_{i}$ using the indicator function:
-$I_{\mathcal{X}_{i}}(x_{i})=\left\{\begin{array}[]{ll}0&\textrm{if }x_{i}\in\mathcal{X}_{i},\\ +\infty&\textrm{otherwise.}\end{array}\right.$ (1.2)
-Then the problem (1.1) can be rewritten as
 $\begin{split}\min_{x_{1},\cdots,x_{N}}&\sum_{i=1}^{N}f_{i}(x_{i})\\ \textrm{subject to}~&\sum_{i=1}^{N}A_{i}x_{i}=c,\end{split}$ (1.1)
-where $f_{i}(x_{i})=\tilde{f}_{i}(x_{i})+I_{\mathcal{X}_{i}}(x_{i})$. Thus, we consider the equivalent reformulation (1.1) throughout this paper for ease of presentation.
-The problem (1.1) appears in various applications that contain the image alignment problem in [33], the robust principal component analysis model with noisy and incomplete data in [36], the latent variable Gaussian graphical model selection in [3, 31] and the quadratic discriminant analysis model in [30].
+where $N\geq 2$, $x_{i}\in\mathbb{R}^{n_{i}},~A_{i}\in\mathbb{R}^{m\times n_{i}},~c\in\mathbb{R}^{m}$ and $f_{i}:\mathbb{R}^{n_{i}}\to\left(-\infty,+\infty\right]~(i=1,2,\cdots,N)$.
+The problem (1.1) appears in various applications such as the image alignment problem in [robust_alignment], the robust principal component analysis model with noisy and incomplete data in [recovering_lowrank], the latent variable Gaussian graphical model selection in [latent_variable, node-based_learning] and the quadratic discriminant analysis model in [discriminant_analysis].
 To solve the problem (1.1), various algorithms have been proposed.
-A simple distributed algorithm for solving (1.1) is dual decomposition [11], which is essentially a dual ascent method or dual subgradient method [35] as follows.
-Consider the Lagrangian for problem (1.1):
+A simple distributed algorithm for solving (1.1) is dual decomposition [Generalized_Lagrange], which is essentially a dual ascent method or dual subgradient method [Minimization_methods].
+To describe it, we consider the Lagrangian for problem (1.1):
 $\mathcal{L}(x_{1},\cdots,x_{N},\lambda)=\sum_{i=1}^{N}f_{i}(x_{i})-\bigg\langle\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\bigg\rangle$ (1.2)
-where $\lambda\in\mathbb{R}^{m}$ is the Lagrangian multiplier or the dual variable. The dual decomposition updates as follows: for $k\geq 1$,
+where $\lambda\in\mathbb{R}^{m}$ is the dual variable. Then, the dual decomposition updates as follows: for $k\geq 1$,
 $\left\{\begin{array}[]{ll}\left(x_{1}^{k+1},x_{2}^{k+1},\cdots,x_{N}^{k+1}\right)=\arg\min_{\{x_{i}\}}\mathcal{L}(x_{1},\cdots,x_{N},\lambda^{k}),\\ \lambda^{k+1}=\lambda^{k}-\alpha_{k}\left(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\right),\end{array}\right$.
@@
 and thus they can be carried out in parallel.
-With a suitable choice of $\alpha_{k}$ and certain assumptions, dual decomposition is guaranteed to converge to an optimal solution [35].
-However, dual decomposition exhibits slow convergence in practice, with a convergence rate of $\mathcal{O}(1/\sqrt{k})$ for general convex problems.
-An effective distributed algorithm for (1.1) can be designed based on the alternating direction method of multipliers (ADMM) [13, 15] using the following augmented Lagrangian:
+With a suitable choice of $\alpha_{k}$ and certain assumptions, dual decomposition is guaranteed to converge to an optimal solution [Minimization_methods].
+However, dual decomposition exhibits slow convergence in practice, with a convergence rate of $\mathit{O}(1/\sqrt{k})$ for general convex problems.
+An effective distributed algorithm for (1.1) can be designed based on the alternating direction method of multipliers (ADMM) [A_dual, Sur_l] using the following augmented Lagrangian:
 $\mathcal{L}_{\rho}(x_{1},\cdots,x_{N};\lambda):=\sum_{i=1}^{N}f_{i}(x_{i})-\bigg\langle\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\bigg\rangle+\frac{\rho}{2}\bigg\|\sum_{i=1}^{N}A_{i}x_{i}-c\bigg\|^{2},$
@@
 In the context of ADMM using the augmented Lagrangian, the primal variable $x_{i}$ can be updated using different approaches.
-One common method is the Gauss-Seidel update, in which the primal variable $x_{i}$ is updated sequentially as follows:
+One common method is the Gauss-Seidel update, in which the primal variable $x_{i}$ is updated sequentially using the most recent values.
+This scheme is commonly referred to as the Gauss-Seidel ADMM, and is given by:
 $\begin{split}x_{i}^{k+1}&=\arg\min_{x_{i}}\mathcal{L}_{\rho}(x_{1}^{k+1},\cdots,x_{i-1}^{k+1},x_{i},x_{i+1}^{k},\cdots,x_{N}^{k},\lambda^{k})\\ &=\arg\min_{x_{i}}f_{i}(x_{i})+\frac{\rho}{2}\bigg\|\sum_{j<i}A_{j}x_{j}^{k+1}+A_{i}x_{i}+\sum_{j>i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\bigg\|_{2}^{2}.\end{split}$
-Alternatively, the Jacobi-type update provides a parallel update scheme for the primal variable $x_{i}$ as follows:
+The Gauss-Seidel ADMM was proposed by Wei and Ozdaglar [Distributed_alternating], where the authors established a convergence rate of $\mathit{O}(1/k)$ under the assumption that each objective function $f_{i}$ is convex.
+Alternatively, the Jacobi-type update, commonly referred to as the Jacobi ADMM, provides a parallel update scheme for the primal variable $x_{i}$ as follows:
 $\begin{split}x_{i}^{k+1}&=\arg\min_{x_{i}}\mathcal{L}_{\rho}(x_{1}^{k},\cdots,x_{i-1}^{k},x_{i},x_{i+1}^{k},\cdots,x_{N}^{k},\lambda^{k})\\ &=\arg\min_{x_{i}}f_{i}(x_{i})+\frac{\rho}{2}\bigg\|A_{i}x_{i}+\sum_{j\neq i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\bigg\|_{2}^{2}.\end{split}$
-Since Gauss-Seidel ADMM updates the primal variables $x_{i}$ using the most recent information, the performance of Jacobi ADMM is typically expected to be worse than that of Gauss-Seidel ADMM.
-For instance, the work [19] found an example for which Jacobi ADMM diverges even when $N=2$.
-This issue was addressed by Deng et al.[6] that introduced the Jacobi-Proximal ADMM involving a proximal term $\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}$ for each $x_{i}$-subproblem with positive semi-definite matrix $P_{i}$ and a damping parameter $\gamma>0$ for the update of $\lambda$ to guarantee its convergence.
+This scheme is suitable for parallel computation, and thus particularly effective for large-scale or big-data scenarios for the model (1.1).
+Deng et al.
+[Parallel_multi-block] established an $\mathit{o}(1/k)$ convergence rate for the Jacobi ADMM, assuming that the objective functions $f_{i}$ are convex and that the matrices $A_{i}$ are full column-rank and mutually near-orthogonal.
+However, with the same penalty parameter $\rho$, the Jacobi ADMM, which updates $x_{i}$ in parallel using previous-iterate information, typically exhibits inferior performance compared to the Gauss-Seidel ADMM, which updates blocks sequentially with the latest iterates.
+In particular, the work [On_full] found an example for which Jacobi ADMM diverges even when $N=2$.
+This issue was addressed by He, Xu, and Yuan [On_the_proximal], who proposed the Jacobi-Proximal ADMM.
+To ensure convergence, they introduced one key modification to the Jacobi ADMM:
+  • Proximal term $\frac{s\rho}{2}\|A_{i}(x_{i}-x_{i}^{k})\|^{2}$: A proximal term $\frac{s\rho}{2}\|A_{i}(x_{i}-x_{i}^{k})\|^{2}$ was added to each $x_{i}$-subproblem, where $\rho>0$ is the penalty parameter and $s>0$ is the proximal coefficient that enforces the new iterate to remain sufficiently close to the previous one.
+Deng et al.
+[Parallel_multi-block] later studied the algorithm in a slightly different parameter setting, which is essentially the same as that of He, Xu, and Yuan [On_the_proximal].
+This variant incorporates the following two modifications to the Jacobi ADMM:
+  • Proximal term $\mathbf{\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}}$: A proximal term $\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}$ was added to each $x_{i}$-subproblem, where $P_{i}$ is a positive semi-definite matrix.
+This promotes strong convexity of the $x_{i}$-subproblems, thereby guaranteeing the uniqueness of their solutions even when the original objective functions are not strictly convex.
+  • Damping parameter $\mathbf{\gamma>0}$: The dual update
+    $\lambda^{k+1}=\lambda^{k}-\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg)$ (1.3)
+    was modified to
+    $\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg),$ (1.4)
+    where $\gamma>0$ is a damping parameter used to control the step size and improve the algorithm’s stability.
 The Jacobi-Proximal ADMM is formulated as in Algorithm 1.
 Algorithm 1 Jacobi-Proximal ADMM
-  Initialize: $x_{i}^{0}~{}(i=1,2,\cdots,N)$ and $\lambda^{0}$;
-  for $k=0,1,\cdots$ do
-     Update $x_{i}$ for $i=1,\cdots,N$ in parallel by:
+1:Initialize: $x_{i}^{0}\ (i=1,2,\ldots,N)$ and $\lambda^{0}$;
+2:for $k=0,1,\ldots$ do
+3:  Update $x_{i}$ for $i=1,\ldots,N$ in parallel by:
 $x_{i}^{k+1}=\arg\min_{x_{i}}\left\{f_{i}(x_{i})+\frac{\rho}{2}\Bigg\|A_{i}x_{i}+\sum_{j\neq i}A_{j}x_{j}^{k}-c-\frac{\lambda^{k}}{\rho}\Bigg\|_{2}^{2}+\frac{1}{2}\big\|x_{i}-x_{i}^{k}\big\|_{P_{i}}^{2}\right\};$ (1.5)
-Update $\textrm{Update }\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg{(}\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg{)}.$ (1.6)
-  end for
+$\textrm{Update }\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg).$ (1.6)
+4:end for
 Here we denote $\|x_{i}\|^{2}_{P_{i}}:=x_{i}^{T}P_{i}x_{i}$.
@@
 There are various choices for $P_{i}$ that can simplify the computation of the $x_{i}$-subproblem.
-Here are two widely used formulations for $P_{i}$ in [7], corresponding to the following methods:
-  1. (1)
+Here are two widely used formulations for $P_{i}$ in [On_the_global], corresponding to the following methods:
+  1. 1.
     The standard proximal method, where $P_{i}$ is taken as $P_{i}=\tau_{i}\mathbf{I}$, with $\tau_{i}>0$.
-  2. (2)
+  2. 2.
     The prox-linear method, where $P_{i}$ is taken as $P_{i}=\tau_{i}\mathbf{I}-\rho A_{i}^{T}A_{i}$, with $\tau_{i}>0$.
-The work of Deng et al.[6] established a sublinear convergence result for the convergence of Jacobi-Proximal ADMM when the proximal matrices $P_{i}$ satisfy $P_{i}\succ(\frac{N}{2-\gamma}-1)A_{i}^{T}A_{i}$ for $i=1,2,\ldots,N$ with $\gamma<2$.
+The work of Deng et al.[Parallel_multi-block] established an $\mathit{o}(1/k)$ convergence result for the convergence of the Jacobi-Proximal ADMM when the proximal matrices $P_{i}$ satisfy $P_{i}\succ(\frac{N}{2-\gamma}-1)A_{i}^{T}A_{i}$ for $i=1,2,\ldots,N$ with $\gamma<2$.
 In particular, under specific choices of $P_{i}$, the convergence conditions reduce to $\tau_{i}>\rho(\frac{N}{2-\gamma}-1)\|A_{i}\|^{2}$ for the standard proximal method and $\tau_{i}>\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}$ for the prox-linear method.
-We then extend the analysis to establish the linear convergence of Jacobi-Proximal ADMM.
-It is well-known that ADMM converges linearly when the cost functions exhibit additional structure, such as strong convexity.
-We prove that the Jacobi-Proximal ADMM converges linearly under the assumption that each cost function $f_{j}$ is strongly convex.
+In this work, we prove that the Jacobi-Proximal ADMM converges linearly under the assumption that each cost function $f_{j}$ is strongly convex.
 We derive a condition on the parameters $\rho>0$ and $\gamma>0$ for linear convergence, and we establish a linear convergence rate dependent on these parameters.
 We refer to Section 2 for the details and we provide a comparison of our convergence result with previous work on the problem (1.1) in Table 1.
-Ref. Convexity Regularity Rank Conditions Rate
-Gauss-Seidel ADMM for $N=2$ [7] $f_{1}$: SC $f_{1}$: $L$-smooth $A_{1}$: full column rank $A_{2}$: full row rank linear
-$f_{1},f_{2}$: SC $f_{1}$: $L$-smooth $A_{2}$: full row rank
-$f_{1}$: SC $f_{1},f_{2}$: $L$-smooth $A_{1}$: full column rank
+Ref.
+Convexity
+Regularity
+Rank
+Conditions
+Rate
+Original ADMM
+for $N=2$
+[Local_linear]
+$f_{1}(x_{1})+f_{2}(x_{2}$)
+$\frac{1}{2}x_{1}^{T}Qx_{1}+q^{T}x_{1}+\frac{1}{2}x_{2}^{T}Rx_{2}+r^{T}x_{2}$
+Local error bound
+$B$: full column rank
+$\mathit{O}(\sigma^{k})$(local)
+Original ADMM
+for $N=2$
+[Linear_convergence_of_the]
+$f_{1},f_{2}$: C
+Global error bound
+$A,B$: full column rank
+$\mathit{O}(\sigma^{k})$
+Original ADMM
+for $N=2$
+[Discerning_the]
+Structured polyhedricity assumption
+(see [Discerning_the, Assump. 1.3])
+$\mathit{O}(\sigma^{k})$(local)
+Gauss-Seidel
+ADMM
+for $N\geq 3$
+[On_the_global_linear]
+$f_{2},\ldots,f_{N}$: SC
+$f_{N}$: $L$-smooth
+$A_{N}$: full row rank
+$\mathit{O}(\sigma^{k})$
+$f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth
+$f_{2},\ldots,f_{N}$: SC
+$f_{1},\ldots,f_{N}$: $L$-smooth
+$A_{1}$: full column rank
+Gauss-Seidel
+ADMM
+for $N\geq 3$
+[On_the_linear]
+$f_{i}(x_{i})=g_{i}(E_{i}x_{i})+h_{i}(x_{i})$
+$g_{i}$: strictly C
+$h_{i}$: C
+$g_{i}$: $L$-smooth
+$\operatorname{epi}(h_{i})$: polyhedral
+$X_{i}$: compact polyhedral sets
+Primal/Dual error bound
+$E_{i}$: full column rank
+$\mathit{O}(\sigma^{k})$
+Jacobi
+ADMM
+for $N\geq 3$
+[Parallel_multi-block]
+$f_{1},\ldots,f_{N}$: C
+$A_{i}$: mutually near-orthogonal
+$\mathit{o}(1/k)$
+Jacobi-Proximal
+ADMM
+for $N\geq 3$
+[Parallel_multi-block]
+$f_{1},\ldots,f_{N}$: C
+$\mathit{o}(1/k)$
+Jacobi-Proximal
+ADMM
+for $N\geq 3$
+This
+work
+$f_{1},\ldots,f_{N}$: SC
+$f_{1},\ldots,f_{N}$: $L$-smooth
+$\big[A_{1}^{T};~A_{2}^{T};~\cdots;~A_{N}^{T}\big]$: full column rank
+$\mathit{O}(\sigma^{k})$
+Table 1: This table summarizes the convergence results for ADMM.
+In the ‘Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex).
+In the ‘Rate’ column, $\mathit{O}(\sigma^{k})$ denotes linear convergence with $0<\sigma<1$ ($\mathit{O}(\sigma^{k})$(local) indicates local linear convergence, and $\mathit{o}(1/k)$ indicates convergence strictly faster than sublinear rate.
+NN denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
+Ref.
+Convexity
+Regularity
+Rank
+Conditions
+Rate
+Proximal
+ADMM
+[Partial_error]
+$f_{1},f_{2}$:C
+Strong Conical Hull
+Intersection Property
+If $P_{1}\succeq 0$: $A_{1},A_{2}$: full column rank
+If $P_{1}\succ 0$: $A_{2}$: full column rank
+$\mathit{O}(\sigma^{k})$
+PADMM-FG
+[On_the_global]
+$f_{1}$: SC
+$f_{1}$: $L$-smooth
+$A_{1}$: full column rank
+$A_{2}$: full row rank
+$\mathit{O}(\sigma^{k})$
+$f_{1},f_{2}$: SC
+$f_{1}$: $L$-smooth
+$A_{2}$: full row rank
+$f_{1}$: SC $f_{1},f_{2}$: $L$-smooth
+$A_{1}$: full column rank
 $f_{1},f_{2}$: SC $f_{1},f_{2}$: $L$-smooth
-Gauss-Seidel ADMM for $N\geq 3$ [26] $f_{2},\ldots,f_{N}$: SC $f_{N}$: $L$-smooth $A_{N}$: full row rank linear
-$f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth
-$f_{2},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $A_{1}$: full column rank
-Gauss-Sediel ADMM for $N\geq 3$ [22] $f_{i}(x_{i})=g_{i}(E_{i}x_{i})+h_{i}(x_{i})$ $g_{i}$: strictly C $h_{i}$: C $g_{i}$: $L$-smooth $E_{i}$: full column rank linear
-Jacobi-Proximal ADMM for $N\geq 3$ [6] $f_{1},\ldots,f_{N}$: C $A_{i}$: full column rank sub-linear
-Jacobi-Proximal ADMM for $N\geq 3$ This work $f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}$: full row rank linear
-Table 1.
-This table summarizes the convergence results for ADMM.
-In the ’Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex).
-In the ’Rate’ column, ’linear’ refers to the cases with linear convergence, while ’sublinear’ indicates the case with sublinear convergence.
-$N$ denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
-1.1. Related works
-The Alternating Direction Method of Multipliers (ADMM) [13, 15] is an effective approach that simplifies the original problem by decomposing it into multiple easier-to-solve minimization subproblems.
-Building on the classical Douglas-Rachford operator splitting method [8], Glowinski and Marrocco [16], along with Gabay and Mercier [14], introduced the well-known alternating direction method of multipliers (ADMM) with Gauss-Seidel decomposition for solving (1.1) when $N=2$.
-It utilizes the augmented Lagrangian for (1.1):
+PADMM-FG
+[Discerning_the]
+Structured polyhedricity assumption
+(see [Discerning_the, Assump. 1.3])
+$\mathit{O}(\sigma^{k})$(local)
+Structured subregularity assumption (see [Discerning_the, Assump. 1.4])
+$A_{1}$: full column rank
+$A_{2}$: full row rank
+Relaxed
+ADMM
+[metric_selection]
+$f_{1}$:SC
+$f_{1}$:$L$-smooth
+AA: full row rank
+$\mathit{O}(\sigma^{k})$
+Relaxed
+ADMM
+[Faster_convergence]
+$f_{1}$: SC
+$f_{1}$: $L$-smooth
+$A_{1}$: full row rank
+$\mathit{O}(e^{-\alpha k})$
+$f_{2}$: SC
+$f_{2}$: $L$-smooth
+$A_{2}$: full row rank
+$f_{2}$: SC $f_{1}$: $L$-smooth
+$A_{1}$: full row rank
+$f_{1}$: SC $f_{2}$: $L$-smooth
+$A_{2}$: full row rank
+Over-Relaxed
+ADMM
+[A_general]
+$f_{1}$:SC, $f_{2}$:C
+$f_{1}$: $L$-smooth
+$A_{2}$: full column rank
+$\mathit{O}(\sigma^{k})$
+Linearized
+ADMM
+[Discerning_the]
+Structured polyhedricity assumption
+(see [Discerning_the, Assump. 1.3])
+$\mathit{O}(\sigma^{k})$(local)
+Structured subregularity assumption (see [Discerning_the, Assump. 1.4])
+$A_{2}$: full row rank
+Table 2: This table summarizes the convergence results for several variants of ADMM in the case $N=2$.
+Here, PADMM-FG refers to the proximal ADMM with the Fortin–Glowinski relaxation step.
+In the ‘Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex).
+In the ‘Rate’ column, $\mathit{O}(\sigma^{k})$ with $0<\sigma<1$ and $\mathit{O}(e^{-\alpha k})$ with $\alpha>0$ both denote linear convergence, while $\mathit{O}(\sigma^{k})$(local) indicates local linear convergence.
+NN denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
+1.1 Related works
+The Alternating Direction Method of Multipliers (ADMM) [A_dual, Sur_l] is an efficient approach that facilitates the solution of structured optimization problems by decomposing them into a sequence of simpler subproblems.
+Building on the classical Douglas–Rachford operator splitting method [On_the], Glowinski and Marrocco [Sur_l], along with Gabay and Mercier [A_dual], introduced the well-known Alternating Direction Method of Multipliers (ADMM) with a Gauss–Seidel-type decomposition for solving (1.1) in the case of $N=2$.
+It employs the augmented Lagrangian for (1.1):
 $\mathcal{L}_{\rho}(x_{1},\cdots,x_{N};\lambda):=\sum_{i=1}^{N}f_{i}(x_{i})-\left<\lambda,\sum_{i=1}^{N}A_{i}x_{i}-c\right>+\frac{\rho}{2}\left\|\sum_{i=1}^{N}A_{i}x_{i}-c\right\|^{2}$
-which incorporates a quadratic penalty of the constraints (with a parameter $\rho>0$) into the Lagrangian.
-At each iteration, the augmented Lagrangian is sequentially minimized with respect to $x_{1}$ and $x_{2}$, followed by an update of the dual variable $\lambda$.
+which incorporates a quadratic penalty term of the constraints (with a penalty parameter $\rho>0$) into the Lagrangian.
+At each iteration, the augmented Lagrangian is sequentially minimized with respect to the primal variables $x_{1}$ and $x_{2}$, followed by an update of the dual variable $\lambda$.
 The ADMM has been extensively investigated for solving two-block convex minimization problems (i.e., $N=2$).
-Its global convergence in the case of $N=2$ has been established in [9, 12].
-In a more recent work by Deng and Yin [7], a generalized version of ADMM was introduced, where proximal terms were added to the two subproblems in (2.2), demonstrating linear convergence under certain assumptions on the strong convexity of functions $f_{1}$ and $f_{2}$, as well as the rank conditions on $A_{1}$ and $A_{2}$.
-Additionally, various studies have explored the convergence rate properties of ADMM when $N=2$ (see, e.g., [28, 10, 18, 1, 29, 20, 32, 21]).
-Although the convergence properties of ADMM for $N=2$ are well understood, its behavior for $N\geq 3$ has remained uncertain for a considerable period.
-In the work by Chen et al.
-[4], a counterexample was provided, demonstrating that without additional conditions, the ADMM for $N\geq 3$ may fail to converge.
+Its global convergence in the case of $N=2$ has been established in [On_the_Douglas, Chapter_ix].
+Han and Yuan [Local_linear] showed that original ADMM achieves a local linear convergence rate for quadratic programs by using an error bound established in [Error_bound].
+Based on this error bound framework, Yang and Han [Linear_convergence_of_the] proved that ADMM attains global linear convergence for convex problems with polyhedral constraints by leveraging a global error bound.
+Yuan, Zeng, and Zhang [Discerning_the] showed that original ADMM also enjoys global linear convergence under the structured polyhedrality assumption by employing variational analysis techniques.
+These scenarios are summarized in Table 1.
+Additionally, various studies have explored the convergence rate properties of original ADMM when $N=2$ (see, e.g., [Splitting_algorithms, Local_linear_linear, On_the_O(1/n), Iteration-complexity_of, On_non-ergodic]).
+In addition to these results on the original ADMM, convergence of several variants for the case of $N=2$ has also been studied, such as the Proximal ADMM [Partial_error], Proximal ADMM with the Fortin–Glowinski relaxation step (PADMM-FG) [On_the_global, Discerning_the], Relaxed ADMM [metric_selection, Faster_convergence], Over-Relaxed ADMM [A_general], and Linearized ADMM [Discerning_the].
+The convergence results of these variants are summarized in Table 2.
+While the convergence results of ADMM is well established for the case $N=2$, its behavior for $N\geq 3$ has long remained unresolved; particularly, Chen et al.
+[The_direct] provided a concrete example in which Gauss-Seidel ADMM diverges for $N=3$ for any penalty parameter $\rho$.
 Building on this, we first discuss studies that establishes the convergence conditions for ADMM using the Gauss-Seidel scheme when $N\geq 3$, where the Gauss-Seidel method updates variables sequentially.
-We refer to the works [2], [23], [5], and [27] that achieved global and sublinear convergence results of the 3-block ADMM.
-Han and Yuan [17] established the global convergence of ADMM (2.2) under the assumption that $f_{1},\cdots,f_{N}$ are all strongly convex and $\gamma$ is restricted to a specific range.
-Both Chen, Shen, and You [5] and Lin, Ma, and Zhang [25] relax the conditions compared to [17] by allowing only $N-1$ functions to be strongly convex, with with Chen et al.
-[5] considering the case $N=3$ and Lin et al.
-[25] considering $N\geq 3$.
-Furthermore, Lin, Ma, and Zhang [25] demonstrated that under these conditions, ADMM (2.2) exhibits a sublinear convergence rate in both the ergodic and non-ergodic senses.
-Lin, Ma, and Zhang [24] further presented sufficient conditions to guarantee a sublinear convergence rate for ADMM when $N\geq 3$.
-Lin, Ma, and Zhang [26] proved the global linear convergence of ADMM under the three scenarios.
-The global R-linear convergence of ADMM has been established by Hong and Luo [22].
+We refer to the works [A_convergent], [convergence_analysis], and [Global_convergence] that achieved global and sublinear convergence results of the 3-block ADMM.
+Han and Yuan [A_note] established the global convergence of ADMM under the assumption that $f_{1},\cdots,f_{N}$ are all strongly convex.
+Building on this, both Chen, Shen, and You [convergence_analysis] and Lin, Ma, and Zhang [On_the_sublinear] relax the assumptions in [A_note].
+In particular, [convergence_analysis], for $N=3$, allows one objective function to be not necessarily strongly convex and also relax the range of the penalty parameter, and propose a relaxed ADMM that incorporates an additional computation of optimal step size and prove its global convergence.
+Lin, Ma, and Zhang [On_the_sublinear] demonstrate sublinear convergence with rate $\textit{O}(1/k)$ in the ergodic sense and $\textit{o}(1/k)$ in the non-ergodic sense when one objective function is convex and the remaining N1N-1 are strongly convex, provided that the penalty parameter lies within a certain range.
+Lin, Ma, and Zhang [On_the_global_linear] proved the global linear convergence of Gauss-Seidel ADMM under the three scenarios, which are detailed in Table 1.
+The global R-linear convergence of Gauss-Seidel ADMM has been established by Hong and Luo [On_the_linear]; the corresponding condition is summarized in Table 1.
 Alternatively, Jacobi updates in ADMM have been extensively studied, offering parallelization advantages for multi-block optimization problems.
-The linear convergence of Gauss-Seidel ADMM, Proximal ADMM, and modified Jacobi ADMM, as shown in [22], is established under several assumptions.
-In [6], it was established that the multi-block Jacobi ADMM converges globally at a rate of $o(1/k)$ in two cases: (i) when all coefficient matrices are nearly mutually orthogonal and possess full column rank, or (ii) when proximal terms are introduced into the subproblems.
-In [19], the multi-block Jacobi ADMM was combined with a relaxation step, achieving a worst-case convergence rate of $O(1/k)$ in both ergodic and non-ergodic senses.
-In a recent work by Xiao, Deng, and Xu [38], the authors introduce the Distributed Jacobi-Proximal ADMM algorithm for solving consensus optimization problems.
-1.2. Contribution
+The linear convergence of Proximal ADMM and modified Jacobi ADMM, as well as that of Gauss-Seidel ADMM, which was mentioned earlier, is established under several assumptions in [On_the_linear].
+Deng et al.
+[Parallel_multi-block] showed that the Jacobi ADMM and Jacobi-Proximal ADMM achieve $\mathit{o}(1/k)$ convergence rate under the conditions summarized in Table 1.
+He, Hou, and Yuan [On_full] proposed a variant of multi-block Jacobi ADMM by incorporating an under-relaxation step into the full Jacobian decomposition framework, achieving a worst-case convergence rate of $O(1/k)$ in both ergodic and non-ergodic senses.
+1.2 Contribution
 The main contribution in this paper are as follows.
-We establish the global linear convergence of Jacobi-Proximal ADMM(Algorithm 1) under the assumption that the objective function $f_{i}$ are strictly convex and $L_{i}$-smooth.
-We establish its convergence at a rate of $O(\sigma^{k})$ for the optimality measure defined in (2.4).
+We establish a global linear convergence of Jacobi-Proximal ADMM(Algorithm 1) under the assumption that the objective functions $f_{i}$ are strictly convex and $L_{i}$-smooth.
 Our numerical results on the LCQP model and the optimal resource allocation problem support the linear convergence result of this work.
-1.3. Organization
+1.3 Organization
 We organize the rest of this paper as follows.
-In Section 2, we present a sufficient condition to ensure the convergence of (2.2).
-Then based on the analysis in Section 2, we prove the global convergence of the Algorithm 1 in Section 3.
-In Section 4, we present convergence results by conducting numerical experiments on two examples.
-Finally, we conclude the paper in Section 5.
-2. Statement of the main result
-In this section, we state the main result of this paper concerning the linear convergence of the Jacobi-Proximal ADMM in Algorithm 1.
+In Section 2, we state the assumptions and the main result of this paper.
+Section 3 is devoted to prove the main result.
+In Section 4, we present numerical experiments supporting the linear convergence result.
+2 Statement of the main result
+In this section, we state the main result of this paper concerning the linear convergence of the Jacobi-Proximal ADMM described in Algorithm 1.
 Since both of the objective function and constraints of (1.1) are sums of terms on individual $x_{i}$’s, the problem can be decomposed into $N$ smaller subproblems, which is solved in a parallel and distributed manner.
@@
 where $\lambda\in\mathbb{R}^{m}$ is the dual variable and parameter $\rho>0$.
+Here we let $n=\sum_{i=1}^{N}n_{i}$ and set $A\in\mathbb{R}^{m\times n}$ by $A=\big[A_{1}\cdots A_{N}\big]$.
+For $k\geq 0$, we denote by $\left(x_{1}^{k},\cdots,x_{N}^{k};\lambda^{k}\right)$ the variable generated by Algorithm 1.
+Also, we set $x^{k}\in\mathbb{R}^{n}$ for $k\geq 0$ and $x^{*}\in\mathbb{R}^{n}$ as $x^{k}=\big[x_{1}^{k};\cdots;x_{N}^{k}\big]$ and $x^{*}=\big[x_{1}^{*};\cdots;x_{N}^{*}\big]$.
+With this notation, we have
+$Ax^{k}=\sum_{i=1}^{N}A_{i}x_{i}^{k}.$
+We use $\|z\|$ to denote the standard Euclidean norm of $z$, and $\langle\cdot,\cdot\rangle$ to denote the standard inner product.
+For a positive definite matrix $M\in\mathbb{R}^{l\times l}$, we define the $M$-norm as $\|z\|_{M}:=\sqrt{z^{T}Mz},\quad\forall z\in\mathbb{R}^{l}$.
+If the matrix $M$ is positive semi-definite, then $\|\cdot\|_{M}$ is a semi-norm.
 We state assumptions on the cost functions which will be used throughout the paper.
-Here we let $n=n_{1}+\cdots+n_{N}$ and set $A\in\mathbb{R}^{m\times n}$ by $A=\big{[}A_{1}\cdots A_{N}\big{]}$.
-Also, we set $x^{k}\in\mathbb{R}^{n}$ for $k\geq 0$ and $x^{*}\in\mathbb{R}^{n}$ as $x^{k}=\big{[}x_{1}^{k};~{}\cdots;~{}x_{N}^{k}\big{]}$ and $x^{*}=\big{[}x_{1}^{*};~{}\cdots;~{}x_{N}^{*}\big{]}$.
-We then have
-$Ax^{k}=\sum_{i=1}^{N}A_{i}x_{i}^{k}.$
 Assumption 2.1.
 The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\left(-\infty,+\infty\right]$ are closed proper convex for $1\leq i\leq N$.
+We assume that there exists a minimizer of problem (1.1) satisfying the KKT condition as follows.
 Assumption 2.2.
@@
 Assumption 2.5.
-We assume that the matrix $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}\in\mathbb{R}^{n\times m}$ with $n=n_{1}+\cdots+n_{N}$ has full column rank.
-Here $P_{i}$ is positive semi-definite for each $1\leq i\leq N$.
-For $k\geq 0$, we denote by $\left(x_{1}^{k},\cdots,x_{N}^{k};\lambda^{k}\right)$ the variable generated by the Jacobi-Proximal ADMM.
+We assume that the matrix $\big[A_{1}^{T};~A_{2}^{T};~\cdots;~A_{N}^{T}\big]\in\mathbb{R}^{n\times m}$ with $n=\sum_{i=1}^{N}n_{i}$ has full column rank.
 We state the main result of this paper.
 Theorem 2.6.
-Suppose Assumptions 2.1-2.5 hold. Choose any $s>0$ such that
+Suppose Assumptions 2.1-2.5 hold and assume that $P_{i}$ is a positive semi-definite matrix for each $1\leq i\leq N$. Choose any $s>0$ such that
 $s\Big[\rho^{2}D\|A_{i}\|+\frac{L}{N}\Big]<\frac{\alpha}{2N},\qquad\forall 1\leq i\leq N,$
-and suppose that $20<\gamma<$ 2, $\rho>0$ are chosen so that there exist values $\xi_{i}>0$ such that
-$\left\{\begin{array}[]{ll}\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}\succ 0,\\ \sum_{i=1}^{N}\xi_{i}<2-\gamma.\end{array}\right.$ (2.3)
-Let $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}$. Here $c_{A}>0$ and $\mu_{s}\in(0,1)$ are constants satisfying
-[…]
-where the first inequality ensures that $c_{A}=\sqrt{\min\textrm{eig}\Big{(}\sum_{i=1}^{N}A_{i}A_{i}^{T}\Big{)}}>0$. Then we have
+and suppose that $0<\gamma<2$, $\rho>0$ are chosen so that there exist values $\xi_{i}>0$ such that
+$\left\{\begin{array}[]{ll}\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}\succ 0,\\ \sum_{i=1}^{N}\xi_{i}<2-\gamma.\end{array}\right$.
+(2.3)
+Define $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}\in(0,1)$ for any $c_{A}\in\Big(0,\frac{1}{\sqrt{2\gamma\rho s}}\Big)$, and $\mu_{s}\in(0,1)$ satisfying
+$\left\{\begin{aligned} &c_{A}^{2}I_{m}\preceq\sum_{i=1}^{N}A_{i}A_{i}^{T},\\ &(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}\preceq\mu_{s}(\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-Ls)I_{n_{i}}.\end{aligned}\right$.
+Then we have
 $\phi\big(x^{k+1},\lambda^{k+1}\big)\leq\sigma\phi\big(x^{k},\lambda^{k}\big),$ (2.4)
-where $\phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is defined as $\phi\big{(}x^{k},\lambda^{k}\big{)}=\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}$.
-As a result of (2.4), we have the following result
+where $\phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is defined as
+$\phi\big(x^{k},\lambda^{k}\big)=\frac{1}{2\gamma\rho}\big\|\lambda^{k}-\lambda^{*}\big\|^{2}+\frac{1}{2}\sum_{i=1}^{N}\big\|x_{i}^{k}-x_{i}^{*}\big\|_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}.$
+As a result of (2.4), we have the following linear convergence result
 $\phi\big(x^{k},\lambda^{k}\big)=O(\sigma^{k}).$
@@
 $\tau_{i}-8s\tau_{i}^{2}-\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}>0.$
-3. Convergence
-In this section, we prove Theorem 2.6. For this, we prepare the several lemmas.
+3 Proof of the main result
+In this section, we give the proof of Theorem 2.6. For this, we prepare several lemmas.
 Lemma 3.1.
@@
 We define $c_{A}\geq 0$ by
-$c_{A}=\min_{\lambda\in\mathbb{R}^{m}\backslash\{0\}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}=\min_{\lambda\in S^{m-1}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}.$
+[…]
 In order to prove that $c_{A}>0$, we assume $c_{A}=0$ and proceed to derive a contradiction. Since $S^{m-1}$ is compact, there exists a minimizer $\lambda_{0}\in S^{m-1}$ of (3.1) such that
@@
 $(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}\prec\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}.$
-Thus, we obtain the final result for $\mu_{s}\in(0,1)$, which is sufficiently close to $1$. ∎
+Thus, we obtain the final result for $\mu_{s}\in(0,1)$, which is sufficiently close to 11. ∎
 Lemma 3.7.
-If one chooses some $\xi_{i}>0$ such that $20<\gamma<$ 2, $\rho>0$, and $s>0$ satisfies the following conditions:
+If one chooses some $\xi_{i}>0$ such that $0<\gamma<2$, $\rho>0$, and $s>0$ satisfies the following conditions:
 $\left\{\begin{array}[]{ll}\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}\succ 0,\\ \sum_{i=1}^{N}\xi_{i}<2-\gamma,\end{array}\right$.
@@
 for $\xi_{j}>0~(1\leq j\leq N)$.
-Adding $\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}$, $-8s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}$, and $\frac{2-\gamma}{\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}$ to both sides of (3.15), we obtain
+Adding $\sum_{i=1}^{N}\big\|x_{i}^{k+1}-x_{i}^{k}\big\|_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}$, $8s\sum_{i=1}^{N}\bigg\|\Big(\rho A_{i}^{T}A_{i}+P_{i}\Big)\big(x_{i}^{k+1}-x_{i}^{k}\big)\bigg\|^{2}$, and $\frac{2-\gamma}{\gamma^{2}\rho}\big\|\lambda^{k+1}-\lambda^{k}\big\|^{2}$ to both sides of (3.15), we obtain
 $\begin{split}&\sum_{i=1}^{N}\big\|x_{i}^{k+1}-x_{i}^{k}\big\|_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}-8s\sum_{i=1}^{N}\bigg\|\Big(\rho A_{i}^{T}A_{i}+P_{i}\Big)\big(x_{i}^{k+1}-x_{i}^{k}\big)\bigg\|^{2}\\ &\quad+\frac{2-\gamma}{\gamma^{2}\rho}\big\|\lambda^{k+1}-\lambda^{k}\big\|^{2}+\frac{2}{\gamma}\Big\langle\lambda^{k}-\lambda^{k+1},A\big(x^{k}-x^{k+1}\big)\Big\rangle\\ &\geq\sum_{i=1}^{N}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}_{\rho A_{i}^{T}A_{i}+P_{i}-8s(\rho A_{i}^{T}A_{i}+P_{i})^{T}(\rho A_{i}^{T}A_{i}+P_{i})-\frac{\rho}{\xi_{i}}A_{i}^{T}A_{i}}\\ &+\bigg(\frac{2-\gamma-\sum_{i=1}^{N}\xi_{i}}{\gamma^{2}\rho}\bigg)\|\lambda^{k+1}-\lambda^{k}\|^{2}\end{split}$
@@
 where $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}$. The proof is done. ∎
-4. Numerical experiments
+4 Numerical experiments
 In this section, we present numerical tests of Algorithm 1 supporting the linear convergence property proved in Theorem 2.6.
-We test the Algorithm 1 for the linearly const rained quadratic programming (LCQP) model [34] and the optimal source allocation problem [37].
-4.1. The LCQP Model
+We test the Algorithm 1 for the linearly const rained quadratic programming (LCQP) model [A_faster] and the optimal source allocation problem [Optimal_scaling].
+In Section 4, all numerical experiments were performed using Python 3.10.16 with NumPy 2.1.3, SciPy 1.10.1, and Matplotlib 3.9.2 on a workstation with an Intel Core i9-12900KS CPU (3.40 GHz) and 32.0 GB RAM.
+4.1 The LCQP Model
 Here we consider the following LCQP model:
@@
 We take $n_{1}=\cdots=n_{N}:=n$ in our experiment.
-We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1.
+We conduct experiments with the damping parameter $\gamma\in\{0.1,0.5,1.5,1.9\}$, and the penalty parameter $\rho$, chosen as $\{0.03,1,5,10\}$ for $N=3$ and $\{10^{-5},0.1,5,10\}$ for $N=10$, in Algorithm 1.
 The matrices involved in this problem are randomly generated through the following procedure:
-  1. (1)
+  1. 1.
     Each $A_{i}\in\mathbb{R}^{m\times n_{i}}(i=1,\ldots,N)$ is randomly generated with i.i.d.
 standard normal entries with satisfying the condition that $\big[A_{1}^{T};\ldots;A_{N}^{T}\big]$ has full column rank.
-  2. (2)
+  2. 2.
     The positive semi-definite matrices $P_{i}\in\mathbb{R}^{n_{i}\times n_{i}}(i=1,\ldots,N)$ are generated by sampling $B\in\mathbb{R}^{n_{i}\times n_{i}}$ with i.i.d.
 standard normal entries, and then symmetrizing as $P_{i}=\frac{1}{2}(B+B^{T})$ with checking that the eigenvalues of $P_{i}$ are non-negative.
-  3. (3)
+  3. 3.
     The positive definite matrices $H_{i}\in\mathbb{S}_{+}^{n_{i}}(i=1,\ldots,N)$ are generated by first sampling $B\in\mathbb{R}^{n_{i}\times n_{i}}$ with i.i.d.
 standard normal entries, and then symmetrizing as $H_{i}=\frac{1}{2}(B+B^{T})$ with checking that the eigenvalues of $H_{i}$ are positive.
-  4. (4)
+  4. 4.
     We first generate the optimal solution $x_{i}^{*}\in\mathbb{R}^{n_{i}}$ and $\lambda^{*}\in\mathbb{R}^{m}$ as follows: $x_{i}^{*}=randn(n_{i},1)$ for $i=1,\ldots,N$, and $\lambda^{*}=randn(m,1)$;
-  5. (5)
+  5. 5.
     Next we define $q_{i}\in\mathbb{R}^{n_{i}}(i=1,\cdots,N)$ and $c\in\mathbb{R}^{m}$ as follows: $q_{i}=-H_{i}x_{i}^{*}+A_{i}^{T}\lambda^{*}$ and $c=\sum_{i=1}^{N}A_{i}x_{i}^{*}$ so that $\left({x_{1}^{*}}^{T},\cdots,{x_{N}^{*}}^{T},{\lambda^{*}}^{T}\right)^{T}$ satisfies the KKT conditions.
@@
 We find that the value $\text{dis}(u^{k})$ decreases to zero linearly as expected in Theorem 2.6.
-Figure 1. Experimental results on 3-block LCQP model for fixed $\gamma$ under varying $\rho$.
-Figure 2. Experimental results on 3-block LCQP model for fixed $\rho$ under varying $\gamma$.
-Figure 3. Experimental results on 10-block LCQP model for fixed $\gamma$ under varying $\rho$.
-Figure 4. Experimental results on 10-block LCQP model for fixed $\rho$ under varying $\gamma$.
-4.2. The Optimal Resource Allocation Problem
+As shown in Figure 1, for fixed values of the damping parameter $\gamma\in\{0.1,0.5\}$, the algorithm exhibits the fastest convergence when $\rho=1$.
+This indicates that a smaller penalty parameter does not necessarily lead to faster convergence, emphasizing the importance of proper choice of the penalty parameter $\rho$ for the fast convergence.
+Figure 2 shows the effect of the damping parameter $\gamma$ under fixed $\rho$.
+The convergence rate improves as $\gamma$ increases when $\rho=0.03$, whereas for $\rho=5$, the algorithm exhibits similar convergence rates across all tested values of $\gamma$, indicating limited sensitivity to the damping parameter.
+Figure 1: Experimental results on LCQP model for $N=3$ ($m=100$, $n=40$), with fixed $\gamma$ and varying $\rho$.
+Figure 2: Experimental results on LCQP model for $N=3$ ($m=100$, $n=40$), with fixed $\rho$ and varying $\gamma$.
+Figure 3: Experimental results on LCQP model for $N=10$ ($m=100$, $n=60$), with fixed $\gamma$ and varying $\rho$.
+Figure 4: Experimental results on LCQP model for $N=10$ ($m=100$, $n=60$), with fixed $\rho$ and varying $\gamma$.
+4.2 The Optimal Resource Allocation Problem
 In this example, we consider the following case:
 $\begin{split}&\min\sum_{i=1}^{N}f_{i}(x_{i})\\ &\textrm{s.t. }\sum_{i=1}^{N}x_{i}=0\end{split}$
-where $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is given as $f_{i}(x_{i})=(1/2)a_{i}(x_{i}-c_{i})^{2}+\log[1+exp(b_{i}(x_{i}-d_{i}))]$ with the coefficients $a_{i},b_{i},c_{i},d_{i}$ generated randomly with uniform distributions on $[0,2]$, $[-2,2]$, $[-10,10]$, $[-10,10]$ respectively.
-We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1.
+where $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is given as $f_{i}(x_{i})=(1/2)a_{i}(x_{i}-c_{i})^{2}+\log[1+exp(b_{i}(x_{i}-d_{i}))]$ with the coefficients $a_{i},b_{i},c_{i},d_{i}$ generated randomly with uniform distributions on $[0,2]$, [2,2][-2,2], [10,10][-10,10], [10,10][-10,10] respectively.
+We conduct experiments with the damping parameter $\gamma\in\{0.1,0.5,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.03,1,5,10\}$ in Algorithm 1.
 We update $x_{i}^{k+1}$ in (1.5) using the equation derived from the optimality given as follows:
@@
 In order to study the convergence behavior of the test algorithms, we test two cases: $6$-block and $20$-block.
-The value $u^{*}=\big{(}x_{1}^{*},\ldots,x_{N}^{*};\lambda^{*}\big{)}$ is obtained numerically by running Algorithm 1 $5,000$ iterations.
+The value $u^{*}=\big(x_{1}^{*},\ldots,x_{N}^{*};\lambda^{*}\big)$ is obtained numerically by running Algorithm 1 $4,000$ iterations.
 We employ the same error metric $\text{dis}(u^{k})$ as defined in (4.2) for this experiment.
@@
 We observe that the value $\text{dis}(u^{k})$ converges linearly to zero, as anticipated in Theorem 2.6.
-Figure 5. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
-Figure 6. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
-Figure 7. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
-Figure 8. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
-5. Conclusion
-In this paper, we consider a multi-block Jacobi-Proximal ADMM algorithm for solving linearly coupled optimization problems in a parallel way.
-We established the linear convergence of the algorithm and derived its convergence rate involving more than two variable blocks when the cost functions are strongly convex and Lipschitz continuous.
+Figure 5 illustrates that, for fixed damping parameters $\gamma\in\{0.1,0.5,1.5\}$, the algorithm achieves the fastest convergence when $\rho=1$.
+This suggests that simply choosing a smaller penalty parameter does not guarantee faster convergence, highlighting the need for proper choice of $\rho$ for the fast convergence.
+In Figure 6, the effect of the damping parameter $\gamma$ is investigated at fixed value of $\rho$.
+When $\rho=0.03$, increasing $\gamma$ leads to improved convergence rates.
+In contrast, for $\rho=5$, the convergence rates remain similar regardless of the value of $\gamma$, implying that the algorithm’s performance is relatively insensitive to changes in $\gamma$.
+Figure 5: Experimental results on the optimal resource allocation model for $N=6$, with fixed $\gamma$ and varying $\rho$.
+Figure 6: Experimental results on the optimal resource allocation model for $N=6$, with fixed $\rho$ under varying $\gamma$.
+Figure 7: Experimental results on the optimal resource allocation model for $N=20$, with fixed $\gamma$ under varying $\rho$.
+Figure 8: Experimental results on the optimal resource allocation model for $N=20$, with fixed $\rho$ under varying $\gamma$.
+5 Conclusion
+In this paper, we study the Jacobi-Proximal ADMM algorithm designed to solve the linearly coupled optimization problems in a parallel way.
+We achieved the first result for the linear convergence of the algorithm and derived its convergence rate when the cost functions are strongly convex and Lipschitz continuous.
 This is the first linear convergence result for the Jacobi-Proximal ADMM.
+We also presented numerical experiments supporting the convergence result.
+Declarations
+Funding
+No funding was received for conducting this study.
+Competing interests
+The authors have no relevant financial or non-financial interests to disclose.
+Code availability
+The Python code used for the numerical experiments in Section 4 is publicly available at https://github.com/hyelinchoi/Jacobi_Proximal_ADMM.